# On Safety and Liveness Filtering Using Hamilton-Jacobi Reachability Analysis

### v1

@@ V-B Least Restrictive Safety Filter
$\pi(x,t)=\begin{cases}\pi_{\text{nom}}(x,t)&V(x)>0\\ \pi_{\text{safe}}^{*}(x)&V(x)=0\end{cases}$ (23)
Similar to the liveness case, the least restrictive safety controller follows the nominal controller when the system is not risk of breaching safety, and it switches to $\pi_{\text{safe}}^{*}(x)$ (11) when the system is on the boundary of the BRT. The LR safety filter in (23) can be obtained using our general framework in (22) by using the following $\widetilde{\mathcal{U}}$:
$\widetilde{\mathcal{U}}(x,t)=\begin{cases}\mathcal{U}&\quad\text{ if }V(x)>0\\ \{\pi_{\text{safe}}^{*}(x)\}&\quad\text{ if }V(x)=0\end{cases}$ (24)
Running Example (Safety). We next demonstrate the least restrictive safety filter on our running example. The corresponding system trajectory and control profiles (for $u_{1}$) are shown in Fig. 7 (purple) and Fig. 8 respectively. Applying the LR filter on top of the baseline MPC controller allows the system to avoid collision with obstacles, but it comes with the shortcoming of an overly aggressive control profile at the boundary of the unsafe set near the obstacles, as the controller switching and the bang-bang nature of $\pi_{\text{safe}}^{*}$ results in a jittery control profile that jumps to the limits of the control authority.
Figure 7: Trajectories for the nominal and the proposed safety filters.
Figure 8: Control profiles for $u_{1}$ (z-velocity) for the nominal controller (gray dashed line) and the proposed safety filters.
@@ V-C Smooth Least Restrictive Safety Filter
$\pi(x,t)=\begin{cases}\pi_{\text{nom}}(x,t)&V(x)>0\\ \pi^{+}(x,t)&V(x)=0\end{cases}$ (25)
where $\pi^{+}(x)$ is obtained by solving the following optimization problem:
$\begin{split}\min_{u\in\mathcal{U}}&\quad||u-\pi_{\text{nom}}(x,t)||_{2}^{2}\\ s.t.&\quad\nabla V(x)\cdot f(x,u,d^{*})=0\end{split}$ (26)
@@
Compared to the least restrictive filter, the smooth least restrictive filter allows the system to take a control action that is “closest” to the nominal control, yet keeping it safe at all times, leading to a smoother control profile. Note that we have dropped the $D_{\tau}$ term from the constraint in (26) because we assume the value function has converged.
@@ V-D Smooth Blending of Performance and Safety
$\begin{split}\min_{u\in\mathcal{U}}&\quad h(u,\pi_{\text{nom}}(x))=||u-\pi_{\text{nom}}(x,t)||_{2}^{2}\\ s.t.&\quad\nabla V(x)\cdot f(x,u,d^{*})\geq-\gamma V(x)\end{split}$ (27)
@@
Similar to the least restrictive and smooth least restrictive filters, the smooth blending filter allows the system state to move towards the boundary of $\mathcal{G}_{\text{unsafe}}$ (i.e., the value function can decrease), but unlike the aforementioned controllers, it limits the rate (determined by the user-defined coefficient $\gamma$ and the value of the current state $x$) at which the value function is allowed to decrease. This phenomenon (a) avoids a sudden switching to $\pi_{\text{safe}}^{*}(x)$, leading to a smoother profile; and, (b) encourages the system to maintain a non-zero distance from the BRT boundary at all times.
@@
The smooth blending filter can be obtained by using $\widetilde{\mathcal{U}}(x,t)=\{u\in\mathcal{U}:\nabla V(x)\cdot f(x,u,d^{*})\geq-\gamma V(x)\}$ within the proposed general safety filter framework. Similar to Lemma 3, it can be shown that $\widetilde{\mathcal{U}}(x,t)\subseteq\mathcal{U}_{\text{safe}}(x)$ for all $\gamma\geq 0$.
@@
Running Example (Safety). Resulting state trajectories of applying the smooth blending filter for two different values of $\gamma$ are shown in blue and light blue in Fig. 7, and their control profiles are shown in Fig. 8. As evident from the control profiles, the smooth blending controllers do not produce any sudden switching since they are free to select controls closer to the nominal controls as long as the decrease in value is bounded by the prescribed rate. This rate is adjusted through the parameter $\gamma$, which determines how “confident” the system feels moving toward the unsafe states at a given value level. For this example we used values of $\gamma=4$ and $\gamma=0.75$. A larger value of $\gamma$ allows the system trajectory to move closer to the obstacles (increasingly behaving like a smooth LR filter), while a smaller value results in a more conservative behavior (behaving similar to the default safety controller). This can also be seen from the system trajectory corresponding to $\gamma=0.75$, wherein the safety filter is pushing the system away from the obstacles to ensure that the value function doesn’t decrease too much (similar to the default safety controller), resulting in a very curvy and inefficient trajectory to the goal. In contrast, the trajectory corresponding to $\gamma=4$ ventures fairly close to the BRT boundary. These trajectories highlight a critical characteristic of smooth blending filter – the system performance can heavily depend on $\gamma$. Thus, a proper tuning of $\gamma$ might be needed to obtain a satisfactory behavior.
@@ V-E Comparison of Different Controllers
Figure 9: Comparison between the baseline MPC controller and proposed safety filters.
@@
The least restrictive filter maintains safety and only gets minimum penetration over obstacles; this slight penetration is due to the switching happening not strictly at $V(x)=0$ but when the value function is negative due to finite grid and time step resolution. The safety enforcement increases the total cost over the trajectory compared to the base MPC controller, because when safety is at risk, it steers the system away from unsafe states which conflicts with the baseline controller objective. The least restrictive filter is the fastest to implement among all safety filters as the default safety controller can be computed quickly using the BRT, without requiring to solve any optimization problem, unlike other safety filters. The downside lies in high energy and high jerk in the controls due to the bang-bang nature of the default safety controller.
The smooth least restrictive filter also suffers from some penetration over obstacles because it uses the same switching condition as the regular least restrictive filter but presents improvements both in the energy and jerk of the controls as it no longer relies on purely bang-bang safety controller, with the drawback of increased computation time per step as (26) needs to be solved each time that safety is at risk.
The final two controllers correspond to smooth blending of performance and safety for two $\gamma$ values. Inclusion of $\gamma$ in the safety filter allows the system to reason about safety before reaching the boundary of the unsafe set, allowing the system to maintain a positive signed distance to obstacles over the whole trajectory. This comes at the expense of an increased mean calculation time, especially for the case with smaller $\gamma$ as the more conservative constraint in (27) is slightly harder to solve. It is also worth noting that a high $\gamma$ behaves quite similarly to the smooth LR filter, without its downside of accidentally penetrating the obstacle due to a last-minute switching.
Overall recommendation. Similar to the liveness case, we overall recommend to use the smooth blending safety filter with a high value of $\gamma$, as it nicely tradeoff performance and safety, while maintaining a low jerk. At the same time, using a very high $\gamma$ can bypass the need for tuning of $\gamma$.
@@ VI Hardware Experiments
We model the dynamics of the system as a Dubins car with $V=0.3~{}m/s$ and $u\in[-0.75,0.75]~{}rad/s$ where the state $[x,y,\varphi]^{T}$ represents the $x$ and $y$ positions and the heading of the robot:
$\dot{x}=V\sin\varphi;\quad\dot{y}=V\cos\varphi;\quad\dot{\varphi}=u,$ (28)
@@
We assume perfect state measurement during our experiments. The system state is obtained through the robot’s internal pose estimation capability given by a combination of measurements from an IMU, optical floor tracking sensor, and wheel encoders. To avoid accidental breach of safety in least restrictive and smooth least restrictive filters due to limited grid resolution and discrete timestep (see Sec. V-B), we use a non-zero threshold to activate the safety controller. Specifically, we use the condition $V(x)=\epsilon$ (instead of $V(x)=0$) to trigger switching to $\pi_{\text{safe}}^{*}$. A value of $\epsilon=0.05$ was used in the experimental results shown.
@@
Least Restrictive Filter: Even though the least restrictive filter keeps the system safe with fast queries of the default safety controller, the bang-bang nature of the safety controller forces the system to take sharp turns. Additionally, it results in high energy control inputs. Together, these results in a high accumulated cost over the robot trajectory.
Smooth Least Restrictive Filter: Similarly to the previously presented simulation results, this method shows improvements compared to the least restrictive filter, both in the total energy and jerk of the control profile. It also keeps lower accumulated cost as the safe control is trying to align with the underlying nominal controller. The drawback comes in the form of an increased computation time per step, as (26) must be solved each time the system safety is at risk. It is also worth noting that using a non-zero threshold to active the QP-based safety controller in (25) avoids any accidental collision with the obstacles, unlike our simulation results.
@@ VII-A Effect of a finite simulation timestep
The filters designed in this paper are based on the assumption that the control input can be applied to the system in a continuous time fashion. However, controller implementation on a real system often involves a zero-order hold, leading to a non-zero simulation timestep $\delta$. If $\delta$ is too large, the proposed controllers may no longer be able to ensure liveness/safety for the underlying continuous time system. For example, an least restrictive filter ensures liveness by switching to $\pi_{\text{live}}^{*}$ at the boundary of the BRT. However, under a finite $\delta$,$\pi_{\text{live}}^{*}$ may no longer be able to keep the system state within the BRT, as the liveness is only guaranteed under a continuous time, state feedback reachability controller. This discrepancy between the continuous-discrete structure is demonstrated in Fig. 12, where we show the trajectory variations generated by the least restrictive filter for different values of $\delta$.
@@
Such effects are more pronounced in the least restrictive filters since there, the system evolves along the BRT boundary. If our simulation timestep is too high, it is highly possible that our system temporarily escape the BRT and enter the unsafe region. One possible solution could be a faster feedback with gradually decreasing $\delta$ as the system approaches the BRT or adaptive $\delta$ depending on the value function. One can also use a discrete-time variant of the BRT computation itself. Alternatively, instead of the switching at the BRT boundary ($V(x)=0$) as seen in Eqn. (23), we can choose to switch with some $\epsilon$ margin from the boundary. This slightly modifies the least restrictive controller as follows,
$\pi(x,t)=\begin{cases}\pi_{\text{nom}}(x,t)&\quad\text{ if }V(x,t)<\epsilon\\ \pi_{\text{live}}^{*}(x,t)&\quad\text{ if }V(x)\geq\epsilon\end{cases}$ (29)
An example corresponding to $\epsilon=-0.1,\delta=0.0005$ is shown with the green line in Fig. 12. As evident from the figure, this strategy is able to maintain liveness at all times and eventually reaches the target set (unlike the pink trajectory using the same $\delta=0.0005$, but $\epsilon=0$ ). This strategy was also employed during our hardware experiments to avoid accidental collisions with the obstacles. Finally, one can use smooth blending filter with a high value of $\gamma$ to avoid liveness/safety loss at the BRT boundary. Nevertheless, a thorough analysis of effect of a finite simulation timstep on system liveness/safety would be a promising future research direction.
@@ VII-B Non-differentiable value functions
The proposed controllers in this paper rely on differentiability of the value function, which may not always be the case [12, 19]. To discuss the impact of non-differentiability of the value function on the proposed controllers, consider the Dubins car dynamics in (28). Moreover, we consider the smooth least restrictive safety filter for our illustration. This requires solving an optimization problem where we find the closest control to the nominal control, such that the condition $\langle\nabla V(x),f(x,u,d^{*})\rangle=0$ with $u\in\mathcal{U}$ is met.
The goal region and the obstacle are shown in light pink and gray respectively in Fig. 13. The nominal planner, unaware of the environment obstacle, plans a straight path to the goal region (blue trajectory).
Correspondingly, the safety controller engages when the system state reaches the BRT boundary (obstacle and BRT shown in top right inlet and boundary point in purple). The value function has a kink at this state, wherein the system has two equally viable option (hard left-turn or hard right-turn) to barely avoid the obstacle. However, at this state, the value function gradient in the theta direction does not exist, as is evident from the bottom right inlet in Fig. 13.
Figure 13: Impact of non-differentiability of the value function on the safety filter. We show example trajectories, the slice for $\theta=\pi/2$, as well as how the value function changes as a function of $\theta$ in the neighborhood of the problematic state (purple dot). $\frac{\partial V}{\partial\theta}$ does not exist in this case.
Thus, one can no longer use the safety filters that rely on the value function gradient (such as a smooth least restrictive filter or a smooth blending filter). One approach to avoid this issue could be to use one-sided gradient of the value function to construct the safety filter. The result of using a left-sided or a right-sided approximation of the gradient is shown in green trajectories in Fig.13. Both approximations are able to keep the system safe while steering it to its goal location. Another solution could be to use (approximate) value functions that are differentiable by construction [5, 15]. However, such approximate BRTs are often conservative in nature; thus, there might be a performance trade-off, depending on the level of conservatism.

### v2

@@ V-B Least Restrictive Safety Filter
$\pi(\mathbf{x},t)=\begin{cases}\pi_{\text{nom}}(\mathbf{x},t)&V(\mathbf{x})>0\\ \pi_{\text{safe}}^{*}(\mathbf{x})&V(\mathbf{x})=0\end{cases}$ (24)
Similar to the liveness case, the least restrictive safety controller follows the nominal controller when the system is not risk of breaching safety, and it switches to $\pi_{\text{safe}}^{*}(\mathbf{x})$ (20) when the system is on the boundary of the BRT. The LR safety filter in (24) can be obtained using our general framework in (23) by using the following $\widetilde{\mathcal{U}}$:
$\widetilde{\mathcal{U}}(\mathbf{x},t)=\begin{cases}\mathcal{U}&\quad\text{ if }V(\mathbf{x})>0\\ \{\pi_{\text{safe}}^{*}(\mathbf{x})\}&\quad\text{ if }V(\mathbf{x})=0\end{cases}$ (25)
Running Example (Safety). For this running example we use an MPC-based nominal controller, whose objective is to minimize distance between the system and the target area, subjected to the system dynamics and the control bounds. To highlight the impacts of safety filtering, the nominal controller is built such that it does not consider obstacle avoidance in its objective. The system state trajectories and control profiles under the nominal controller are shown in grey in Fig. 7 and Fig. 8 respectively. As expected, the nominal controller fails to satisfy the safety constraint (i.e. avoiding the obstacles).
We next demonstrate the least restrictive safety filter acting over the nominal controller. The corresponding system trajectory and control profiles (for $u_{1}$) are shown in Fig. 7 (purple) and Fig. 8 respectively. This filtering technique allows the system to avoid collision with obstacles, but it comes with the shortcoming of an overly aggressive control profile at the boundary of the unsafe set near the obstacles, as the controller switching and the bang-bang nature of $\pi_{\text{safe}}^{*}$ results in a jittery control profile that jumps to the limits of the control authority.
Figure 8: Control profiles for $u_{1}$ (z-velocity) for the nominal controller (gray dashed line), the proposed safety filters, and a CBF-based filter.
@@ V-C Smooth Least Restrictive Safety Filter
$\pi(\mathbf{x},t)=\begin{cases}\pi_{\text{nom}}(\mathbf{x},t)&V(\mathbf{x})>0\\ \pi^{+}(\mathbf{x},t)&V(\mathbf{x})=0\end{cases}$ (26)
where $\pi^{+}(\mathbf{x})$ is obtained by solving the following optimization problem:
$\begin{split}\min_{u\in\mathcal{U}}&\quad||u-\pi_{\text{nom}}(\mathbf{x},t)||_{2}^{2}\\ s.t.&\quad\nabla V(\mathbf{x})\cdot f(\mathbf{x},u,d^{*})=0\end{split}$ (27)
@@
Compared to the least restrictive filter, the smooth least restrictive filter allows the system to take a control action that is “closest” to the nominal control, yet keeping it safe at all times, leading to a smoother control profile. Note that we have dropped the $D_{t}$ term from the constraint in (27) because we assume the usage of the converged value function.
@@ V-D Smooth Blending of Performance and Safety
$\begin{split}\min_{u\in\mathcal{U}}&\quad h(u,\pi_{\text{nom}}(\mathbf{x},t))=||u-\pi_{\text{nom}}(\mathbf{x},t)||_{2}^{2}\\ s.t.&\quad\nabla V(\mathbf{x})\cdot f(\mathbf{x},u,d^{*})\geq-\gamma V(\mathbf{x})\end{split}$ (28)
@@
Similar to the least restrictive and smooth least restrictive filters, the smooth blending filter allows the system state to move towards the boundary of $\mathcal{G}_{\text{unsafe}}$ (i.e., the value function can decrease), but unlike the aforementioned controllers, it limits the rate (determined by the user-defined coefficient $\gamma$ and the value of the current state $\mathbf{x}$) at which the value function is allowed to decrease. This phenomenon (a) avoids a sudden switching to $\pi_{\text{safe}}^{*}(\mathbf{x})$, leading to a smoother profile; and, (b) encourages the system to maintain a non-zero distance from the BRT boundary at all times.
@@
The smooth blending filter can be obtained by using $\widetilde{\mathcal{U}}(\mathbf{x},t)=\{u\in\mathcal{U}:\nabla V(\mathbf{x})\cdot f(\mathbf{x},u,d^{*})\geq-\gamma V(\mathbf{x})\}$ within the proposed general safety filter framework. Similar to Lemma 3, it can be shown that $\widetilde{\mathcal{U}}(\mathbf{x},t)\subseteq\mathcal{U}_{\text{safe}}(\mathbf{x})$ for all $\gamma\geq 0$.
@@
Running Example (Safety). Resulting state trajectories of applying the smooth blending filter for two different values of $\gamma$ are shown in blue and light blue in Fig. 7, and their control profiles are shown in Fig. 8. As evident from the control profiles, the smooth blending controllers do not produce any sudden switching since they are free to select controls closer to the nominal controls as long as the decrease in value is bounded by the prescribed rate. This rate is adjusted through the parameter $\gamma$, which determines how “confident” the system feels moving toward the unsafe states at a given value level. For this example we used values of $\gamma=4$ and $\gamma=0.75$. A larger value of $\gamma$ allows the system trajectory to move closer to the obstacles (increasingly behaving like a smooth LR filter), while a smaller value results in a more conservative behavior (behaving similar to the default safety controller). This can also be seen from the system trajectory corresponding to $\gamma=0.75$, wherein the safety filter is pushing the system away from the obstacles to ensure that the value function doesn’t decrease too much (similar to the default safety controller), resulting in a very curvy and inefficient trajectory to the goal. In contrast, the trajectory corresponding to $\gamma=4$ ventures fairly close to the BRT boundary. These trajectories highlight a critical characteristic of smooth blending filter – the system performance can heavily depend on $\gamma$. Thus, a proper tuning of $\gamma$ is needed to obtain a satisfactory behavior.
Even though the proposed filter ensures safety for any value of $\gamma\geq 0$, we do not study mechanisms to compute an appropriate $\gamma$ in this work. An appropriate value of $\gamma$ will depend on how the user characterizes performance, safety conservativeness, and the compromise between them. The choice of gamma further depends on the system dynamics, failure set, control and disturbance bounds, and the nominal controller. For our running example, we simulated various values of $\gamma$ and picked two values to highlight the extreme behaviors that $\gamma$ can cause in the filtered trajectory.
@@ V-E Comparison of Different Controllers
Figure 9: Comparison between the baseline MPC controller, proposed safety filters and a CBF-based filter.
@@
Considering the smooth blending filter’s CBF-inspired nature, we also implement a CBF-based safety filter applied to the same baseline MPC controller. Since the dynamical system defined by (21) is control-affine and the relative degree of $l(\mathbf{x})$ to the dynamics is one, we can use the approach presented in [27] to use it as a High Order CBF. We choose the hyperparameters for this approach such that the CBF covers $\mathcal{G}_{\text{unsafe}}$, and the hyperparameter that corresponds to our parameter $\gamma$ is set to $4$ for a fair comparison.
The least restrictive filter maintains safety and only gets minimum penetration over obstacles; this slight penetration is due to the switching happening not strictly at $V(\mathbf{x})=0$ but when the value function is negative due to finite grid and time step resolution. The safety enforcement increases the total cost over the trajectory compared to the base MPC controller, because when safety is at risk, it steers the system away from unsafe states which conflicts with the baseline controller objective. The least restrictive filter is the fastest to implement among all safety filters as the default safety controller can be computed quickly using the BRT, without requiring to solve any optimization problem, unlike other safety filters. The downside lies in high energy and high jerk in the controls due to the bang-bang nature of the default safety controller.
The smooth least restrictive filter also suffers from some penetration over obstacles because it uses the same switching condition as the regular least restrictive filter, but presents improvements both in the energy and jerk of the controls as it no longer relies on purely bang-bang safety controller, with the drawback of increased computation time per step as (27) needs to be solved each time that safety is at risk.
The final three controllers correspond to smooth blending of performance and safety for two $\gamma$ values and the CBF-based filter. Inclusion of the $-\gamma V(\mathbf{x})$ term in the safety filter allows the system to reason about safety before reaching the boundary of the unsafe set, allowing the system to maintain a positive signed distance to obstacles over the whole trajectory. This comes at the expense of an increased mean calculation time, especially for the case with smaller $\gamma$ as the more conservative constraint in (28) is slightly harder to solve. The characteristics mentioned so far are shared with the CBF-based filter, where we observe very similar results to that of a smooth blending filter with $\gamma=4$. This is expected since we are essentially solving an equivalent optimization problem. We also observe some increased cost for the CBF-based filter, as it leads to a more conservative unsafe set compared to the BRT, which corresponds to the minimal unsafe set. The behavior of the CBF-based safety filter becomes even more conservative or overly optimistic if the hyperparameter during the CBF synthesis is not properly tuned. Thus, the safety and conservativeness of a CBF-based filter are directly affected by its synthesis, which can be challenging in practice for general nonlinear systems. It is also worth noting that a high $\gamma$ behaves quite similarly to the smooth LR filter, without its downside of accidentally penetrating the obstacle due to a last-minute switching.
Overall recommendation. Similar to the liveness case, we recommend to use the smooth blending safety filter with a high value of $\gamma$, as it nicely tradeoff performance and safety, while maintaining a low jerk. At the same time, using a very high $\gamma$ can bypass the need for tuning it.
@@ VI Hardware Experiments
We model the dynamics of the system as a Dubins car with $V=0.3~{}m/s$ and $u\in[-0.75,0.75]~{}rad/s$ where the state $\mathbf{x}=[x,y,\varphi]^{T}$ represents the $x$ and $y$ positions and the heading of the robot:
$\dot{x}=V\sin\varphi;\quad\dot{y}=V\cos\varphi;\quad\dot{\varphi}=u$ (29)
@@
We assume perfect state measurement during our experiments. The system state is obtained through the robot’s internal pose estimation capability given by a combination of measurements from an IMU, optical floor tracking sensor, and wheel encoders. To avoid accidental breach of safety in least restrictive and smooth least restrictive filters due to limited grid resolution and discrete timestep (see Sec. V-B), we use a non-zero threshold to activate the safety controller. Specifically, we use the condition $V(\mathbf{x})=\epsilon$ (instead of $V(\mathbf{x})=0$) to trigger switching to $\pi_{\text{safe}}^{*}$. A value of $\epsilon=0.05$ was used in the experimental results shown.
@@
Least Restrictive Filter: Even though the least restrictive filter keeps the system safe with fast queries of the default safety controller, the bang-bang nature of the safety controller forces the system to take the sharpest turn possible as it only reasons about maximally increasing safety. Additionally, it results in high energy control inputs. Together, these results in a high accumulated cost over the robot trajectory.
Smooth Least Restrictive Filter: Similarly to the previously presented simulation results, this method shows improvements compared to the least restrictive filter, both in the total energy and jerk of the control profile. It also keeps lower accumulated cost as the safe control is trying to align with the underlying nominal controller. The drawback comes in the form of an increased computation time per step, as (27) must be solved each time the system safety is at risk. It is also worth noting that using a non-zero threshold to activate the QP-based safety controller in (26) avoids any accidental collision with the obstacles, unlike our simulation results.
@@ VII-A Effect of a finite simulation timestep
The filters designed in this paper are based on the assumption that the control input can be applied to the system in a continuous time fashion. However, controller implementation on a real system often involves a zero-order hold, leading to a non-zero simulation timestep $\delta$. If $\delta$ is too large, the proposed controllers may no longer be able to ensure liveness/safety for the underlying continuous time system. For example, a least restrictive filter ensures liveness by switching to $\pi_{\text{live}}^{*}$ at the boundary of the BRT. However, under a finite $\delta$, $\pi_{\text{live}}^{*}$ may no longer be able to keep the system state within the BRT, as the liveness is only guaranteed under a continuous time, state feedback reachability controller. This discrepancy between the continuous-discrete structure is demonstrated in Fig. 12, where we show the trajectory variations generated by the least restrictive filter for different values of $\delta$.
@@
Such effects are more pronounced in the least restrictive filters since there, the system evolves along the BRT boundary. If our simulation timestep is too high, it is highly possible that our system temporarily escape the BRT and enter the unsafe region. One possible solution could be a faster feedback with gradually decreasing $\delta$ as the system approaches the BRT or adaptive $\delta$ depending on the value function. One can also use a discrete-time variant of the BRT computation itself. Alternatively, instead of the switching at the BRT boundary ($V(\mathbf{x})=0$) as seen in Eqn. (24), we can choose to switch with some $\epsilon$ margin from the boundary. This slightly modifies the least restrictive controller as follows,
$\pi(\mathbf{x},t)=\begin{cases}\pi_{\text{nom}}(\mathbf{x},t)&\quad\text{ if }V(\mathbf{x},t)<\epsilon\\ \pi_{\text{live}}^{*}(\mathbf{x},t)&\quad\text{ if }V(\mathbf{x})\geq\epsilon\end{cases}$ (30)
An example corresponding to $\epsilon=-0.1,\delta=0.0005$ is shown with the green line in Fig. 12. As evident from the figure, this strategy is able to maintain liveness at all times and eventually reaches the target set (unlike the pink trajectory using the same $\delta=0.0005$, but $\epsilon=0$). This strategy was also employed during our hardware experiments to avoid accidental collisions with the obstacles. Nevertheless, a thorough analysis of the effect of a finite simulation timestep on system liveness/safety would be a promising future research direction.
@@ VII-B Non-differentiable value functions
The proposed controllers in this paper rely on differentiability of the value function, which may not always be the case [13, 20]. To discuss the impact of non-differentiability of the value function on the proposed controllers, consider the Dubins car dynamics in (29). Moreover, we consider the smooth least restrictive safety filter for our illustration. This requires solving an optimization problem where we find the closest control to the nominal control, such that the condition $\langle\nabla V(\mathbf{x}),f(\mathbf{x},u,d^{*})\rangle=0$ with $u\in\mathcal{U}$ is met.
First we consider the scenario where the value function is being numerically calculated using a grid based approach [9],[21]. The goal region and the obstacle are shown in light pink and gray respectively in Fig. 13(left). The nominal planner, unaware of the environment obstacle, plans a straight path to the goal region (blue trajectory).
Correspondingly, the safety controller engages when the system state reaches the BRT boundary (purple point). The value function has a kink at this state, wherein the system has two equally viable option (hard left-turn or hard right-turn) to barely avoid the obstacle. However, at this state, the value function gradient in the theta direction does not exist, as is evident from the inlet in Fig. 13(left) showing the variations on value around this point as a function of $\theta$
Figure 13: (Left) BRT slice for $\theta=\pi/2$. The value function computed using a grid-based approach exhibits non-differentiability at the purple point. Inlet: the value function plot at the purple point as a function of theta. The value function exhibits a kink at $\theta=\pi/2$. The green trajectories are computed using one-sided gradients. (Right) A smooth approximation of the value function computed using a learning-based method (the black dashed line is the BRT corresponding to the grid-based approach). The value function now changes smoothly as a function of $\theta$ at the purple point.
Thus, one can no longer use the safety filters that rely on the value function gradient (such as a smooth least restrictive filter or a smooth blending filter). One approach to avoid this issue could be to use one-sided gradient of the value function to construct the safety filter. The result of using a left-sided or a right-sided approximation of the gradient is shown in green trajectories in Fig.13(left). Both approximations are able to keep the system safe while steering it to its goal location.
A second approach is to obtain a conservative approximation of the value function that is differentiable and then construct safety/liveness filters using this approximation. For illustration, here we use a learning-based method, DeepReach, to approximate the value function using neural network representations that are differentiable by construction [4, 16]. In Fig.13(right), we show a comparison of the BRT obtained through DeepReach and the previously discussed grid-based method. The inlet in Fig. 13(right) shows how the learned value function is differentiable at the tip of the BRT, unlike the grid-based value function. Moreover, the BRT obtained through the learning method encompasses the numerically approximated value function, which keeps all the safety guarantees intact at the expense of a slightly more conservative BRT, for completeness the smooth least restrictive filtered trajectory using this differentiable representation is shown in green.